# Adapting to Non-Stationary Environments: Multi-Armed Bandit Enhanced Retrieval-Augmented Generation on Knowledge Graphs

### v1

@@ Abstract
Despite the superior performance of Large language models on many NLP tasks, they still face significant limitations in memorizing extensive world knowledge. Recent studies have demonstrated that leveraging the Retrieval-Augmented Generation (RAG) framework, combined with Knowledge Graphs that encapsulate extensive factual data in a structured format, robustly enhances the reasoning capabilities of LLMs. However, deploying such systems in real-world scenarios presents challenges: the continuous evolution of non-stationary environments may lead to performance degradation and user satisfaction requires a careful balance of performance and responsiveness. To address these challenges, we introduce a Multi-objective Multi-Armed Bandit enhanced RAG framework, supported by multiple retrieval methods with diverse capabilities under rich and evolving retrieval contexts in practice. Within this framework, each retrieval method is treated as a distinct “arm”. The system utilizes real-time user feedback to adapt to dynamic environments, by selecting the appropriate retrieval method based on input queries and the historical multi-objective performance of each arm. Extensive experiments conducted on two benchmark KGQA datasets demonstrate that our method significantly outperforms baseline methods in non-stationary settings while achieving state-of-the-art performance in stationary environments. Code and data are available at https://anonymous.4open.science/r/Dynamic-RAG-78D3
Introduction
@@
Deep Multi-objective Contextual Bandits
@@
5:     Select an retriever (arm) $a$ based on the selection strategy described in Arm Selection Strategy:.
@@
Multi-Objective Optimization with GGI: To balance these objectives, we compute the multi-objective GGI value, which integrates accuracy and efficiency metrics, further detail of GGI property can be found in Generalized Gini Index. The GGI values for each objective are calculated as follows:
@@
Experiment
Dataset & Setup
@@
Implementations: To leverage the strengths of various retrieval methods, our MAB-enhanced RAG system, along with other router-based RAG approaches, utilizes DECAF, ChatKBQA, and Reason-on-Graph as its action space. We consistently use Llama-2-7b-chat-hf (Touvron et al. 2023) as the LLM generator, applying a standard RAG prompt (LlamaIndex 2024) across all methods to ensure a fair comparison. All experiments are conducted on the Nvidia Tesla V100 graphical card with the Intel Xeon Platinum 8255C CPU.
@@
Research Questions and Main Results
@@
Related Work
KG-based RAG Systems: Retrieval-Augmented Generation (RAG)(Lewis et al. 2020) mitigates the hallucination issue of LLMs by retrieving external knowledge to enhance the accuracy and reliability of generation content. Recent RAG advancements have increasingly incorporated Knowledge Graphs (KGs) (Luo et al. 2023c; Sun et al. 2023; Xu et al. 2024; He et al. 2024), which store structured factual information, enabling more systematic reasoning by LLMs (Pan et al. 2024). KGs support diverse retrieval methods, each with different capabilities and costs, as detailed in Retrieval methods on KG.
@@ Conclusion
Appendix
Generalized Gini Index
To further leverage the unique strengths of each method and enhance the overall efficacy of the RAG pipeline. This necessitates the formulation of a multi-objective optimization problem.
The Generalized Gini Index (GGI) emerges as a crucial tool in this context, offering a sophisticated framework for equitably balancing diverse criteria in multi-objective optimization scenarios. The need for such an optimization arises from the varied and often conflicting objectives associated with different retrieval methods. For instance, while one method may excel in accuracy, another might offer benefits in terms of speed. The challenge, therefore, lies in achieving an optimal balance that maximizes the overall performance of the RAG system.
GGI has been characterized by (Weymark 1981). It encodes both efficiency as it is monotone with Pareto dominance and fairness as it is non-increasing with Pigou-Dalton transfers(Jenkins 2017). Informally, a Pigou-Dalton transfer involves increasing a lower-valued objective while decreasing a higher-valued objective by an equivalent amount, without altering the order between the two objectives. This operation seeks to equilibrate the cost vector. Formally, the GGI adheres to the following fairness principle: for any $x_{i}<x_{j}$,
$\forall\epsilon\in(0,x_{j}-x_{i}),\quad G_{w}(\mathbf{x}+\epsilon\mathbf{e}_{i}-\epsilon\mathbf{e}_{j})\leq G_{w}(\mathbf{x})$ (6)
where $\mathbf{e}_{i}$ and $\mathbf{e}_{j}$ are two vectors of the canonical basis. Consequently, among vectors with an equal sum, the optimal cost vector (with respect to GGI) is the one that features uniform values across all objectives, provided that such a distribution is feasible.
MAB enhanced RAG systems with LLM variants
We evaluated our MAB-enhanced RAG system using various LLM generators, including Chatglm3 (Du et al. 2021) and Mistral (Jiang et al. 2023a). The results, depicted in Fig. 5, demonstrate that our system significantly improves RAG performance compared to traditional systems that utilize only a single retriever. This enhancement underscores the robustness and adaptability of our MAB-enhanced RAG approach.
Figure 5: MAB enhanced RAG systems with LLM variants under stationary environments
Case study
We further present a case study Fig. 6 to illustrate the effectiveness of our system, the query inquires about influences on Frank Lloyd Wright. While dense retrieval provides a fast response, it introduces noisy and irrelevant data. The LLM agent retriever, though accessing structured knowledge graphs, fails to deliver accurate information, focusing instead on peripheral details like professions. However, the SPARQL-retriever can give a very accurate query code for this question with clear conditions, and successfully fetch all results that support our systems to offer a nuanced and accurate response.
Figure 6: SPARQL-Based Retriever gives the most accurate context
Figure 7: KG Agent gives the most accurate context
Retrieval methods on KG
Dense retrieval: Dense retrieval methods (Zhang et al. 2023; Karpukhin et al. 2020) standardize and segment diverse document formats like PDF, HTML, Word, and Markdown into plain text, which is then transformed into vector embeddings for efficient searching (Karpukhin et al. 2020). Utilizing a pre-trained language model (Devlin et al. 2018), DPR creates dense embeddings from question-passage pairs, significantly enhancing accuracy over BM25 and ORQA in open Natural Questions. Additionally, recent developments by (Zhang et al. 2023) have tailored embedding models to meet the varied retrieval demands of LLMs with techniques like knowledge distillation and multi-task fine-tuning.
There are two main approaches to applying dense retrieval to Knowledge Graph data: directly searching the textual data (e.g. Wikipedia) that constitutes the KG (Ennen et al. 2023; Karpukhin et al. 2020), and linearizing KGs into text corpora (Yu et al. 2022; Xie et al. 2022), which translates structured knowledge into natural language form.
Dense retrieval methods are typically fast with pre-built vector bases, but it rely on embedding models for query reasoning, often falling short in complex, multi-hop retrieval tasks that demand greater analytical depth (BehnamGhader, Miret, and Reddy 2023).
SPARQL generator: Due to the structured representation and storage, it is efficient to access structured data using query languages (e.g. SPARQL) or specific algorithms (e.g., triple search for knowledge graphs). ChatKBQA (Luo et al. 2023a) proposes generating the logical form with fine-tuned LLMs first, then retrieving and replacing entities and relations through an unsupervised retrieval method, which improves both generation and retrieval more straightforwardly. StructGPT (Jiang et al. 2023b) constructs the specialized interfaces to collect relevant evidence from structured data, and let LLMs concentrate on the reasoning task based on the collected information.
KG agent-based retrieval: Differing from SPARQL generator retrieval, LLM agent-based retrieval adopts a tightly coupled ”LLM-KG” paradigm. In this setup, agents such as LLMs navigate through relations and entities on Knowledge Graphs and construct a reasoning path for answering queries. (Sun et al. 2023) introduces a new approach called Think-on-Graph (ToG), in which the LLM agent iteratively executes beam search on KG, discovers the most promising reasoning paths, and returns the most likely reasoning results. (Luo et al. 2023c) presents a planning retrieval-reasoning framework, where RoG first generates relation paths grounded by KGs as faithful plans. These plans are then used to retrieve valid reasoning paths from the KGs for LLMs to conduct faithful reasoning.
RAG Prompts Used in Experiments
Listing 1: RAG prompt
System prompt:
Based on the context information provided, and not on prior knowledge, please answer the given question.
Context: {context}
Question:
Please answer the query: {query}
Statistics of datasets
Table 4: Statistics of datasets.
Datasets Train Test Max hop
WebQSP 2,826 1,628 2
CWQ 27,639 3,531 4

### v2

@@ Abstract
Despite the superior performance of Large language models on many NLP tasks, they still face significant limitations in memorizing extensive world knowledge. Recent studies have demonstrated that leveraging the Retrieval-Augmented Generation (RAG) framework, combined with Knowledge Graphs that encapsulate extensive factual data in a structured format, robustly enhances the reasoning capabilities of LLMs. However, deploying such systems in real-world scenarios presents challenges: the continuous evolution of non-stationary environments may lead to performance degradation and user satisfaction requires a careful balance of performance and responsiveness. To address these challenges, we introduce a Multi-objective Multi-Armed Bandit enhanced RAG framework, supported by multiple retrieval methods with diverse capabilities under rich and evolving retrieval contexts in practice. Within this framework, each retrieval method is treated as a distinct “arm”. The system utilizes real-time user feedback to adapt to dynamic environments, by selecting the appropriate retrieval method based on input queries and the historical multi-objective performance of each arm. Extensive experiments conducted on two benchmark KGQA datasets demonstrate that our method significantly outperforms baseline methods in non-stationary settings while achieving state-of-the-art performance in stationary environments. Code and data are available at https://github.com/FUTUREEEEEE/Dynamic-RAG
1 Introduction
@@ 2.1 Problem Setup
2.2 Deep Multi-objective Contextual Bandits
@@
5:     Select an retriever (arm) $a$ based on the selection strategy described in Section 2.2.
@@
Multi-Objective Optimization with GGI: To balance these objectives, we compute the multi-objective GGI value, which integrates accuracy and efficiency metrics, further detail of GGI property can be found in Section 6.1. The GGI values for each objective are calculated as follows:
@@
3 Experiment
3.1 Dataset & Setup
@@
Implementations: We consistently use Llama-2-7b-chat-hf (Touvron et al. 2023) as the LLM generator, applying a standard RAG prompt (LlamaIndex 2024) across all methods to ensure a fair comparison. All experiments are conducted on the Nvidia Tesla V100 graphical card with the Intel Xeon Platinum 8255C CPU. See Section 6.3 for detail set up.
@@
3.2 Research Questions and Main Results
@@
4 Related Work
KG-based RAG Systems: Retrieval-Augmented Generation (RAG)(Lewis et al. 2020) mitigates the hallucination issue of LLMs by retrieving external knowledge to enhance the accuracy and reliability of generation content. Recent RAG advancements have increasingly incorporated Knowledge Graphs (KGs) (Luo et al. 2023c; Sun et al. 2023; Xu et al. 2024; He et al. 2024), which store structured factual information, enabling more systematic reasoning by LLMs (Pan et al. 2024). KGs support diverse retrieval methods, each with different capabilities and costs, as detailed in Section 6.6.